OpenStudy (anonymous):

prove that (sinx/(1-cotx)) - (cosx/(tanx-1)) = sinx + cosx

OpenStudy (jdoe0001):

hmm what would be the basic trig identity of cotangent?

OpenStudy (zpupster):

on LHS take first term- chnge it to sin and cos \[\sin x/1-(\frac{ cosx }{ sinx })= \frac{ \sin^2x }{ sinx-cosx }\] now do the same with the 2nd term \[ -\frac{ \cos^2x }{ sinx-cosx }\] can you see it from here??

OpenStudy (anonymous):

No, I can't

OpenStudy (jdoe0001):

hmm what would be the basic trig identities for cot and tan functions?

OpenStudy (zpupster):

ok put terms together \[(\sin^2x-\cos^2x)/sinx-cosx\] see where to go now??

OpenStudy (anonymous):

yes, Thank you!

OpenStudy (zpupster):

good job!!